# GauSSmart: Enhanced 3D Reconstruction through 2D Foundation Models and Geometric Filtering

### v2

@@ 1 Introduction
Three-dimensional (3D) reconstruction has emerged as a prominent area of research since the introduction of Neural Radiance Fields (NeRF) [36]. More recently, Gaussian Splatting [20] has advanced scene reconstruction by exploiting the mathematical properties of 3D Gaussians, which allow for both accurate modeling and efficient rendering. Nevertheless, these methods frequently rely on unstructured point clouds, where uneven density and photogrammetry artifacts often introduce noise, limiting the fidelity of the reconstructed geometry.
@@
Given the spurious conditions of the input point clouds, we first apply a convex outlier removal procedure to eliminate unnecessary points introduced by interpolation errors in COLMAP reconstructions. In addition, we present a framework that leverages 2D foundation models such as DINOv3 [50] and SAM [22] to guide point cloud densification. Specifically, we increase the density of underrepresented segments by sampling new points according to the covariance structure of the original points within each segment mask. Finally, we introduce a novel loss function built on DINOv3 embeddings, which is differentiable and seamlessly integrates into the Gaussian Splatting pipeline [18], enabling improved semantic alignment during optimization.
@@ 2 Related Works
Outlier Removal. Outlier removal is a key pre-processing step for noisy point clouds, with two widely used methods being Statistical Outlier Removal (SOR) and Radius Outlier Removal (ROR), both popularized in the Point Cloud Library [48, 47]. SOR computes the mean distance of each point to its $k$-nearest neighbors and rejects those deviating from the global mean–standard deviation range, while ROR removes points with fewer than a threshold number of neighbors within a fixed radius. These lightweight filters are usually combined with robust model-fitting methods such as RANSAC [13] or Trimmed ICP [7], surface-fitting and projection approaches like Point Set Surfaces [1], graph-based and variational denoisers [67, 4], and learning-based frameworks such as PointCleanNet [44].
@@
Novel View Synthesis Neural Radiance Fields (NeRF)[36] model view-dependent color and density using a multilayer perceptron (MLP) to synthesize novel views. Subsequent variants have improved sampling strategies, scalability, and geometric fidelity[2, 30, 28, 3, 62, 5, 38, 55, 61, 35, 14, 29]. Despite these advances, conventional NeRF models remain computationally expensive and often fail to accurately reconstruct large-scale or texture-poor regions. More recently, scene representations based on projected Gaussian primitives have enabled fast, high-quality rendering while remaining scalable to large scenes [20, 18, 9, 11, 17]. In particular, 2D Gaussian Splatting [18] improves geometric alignment by orienting primitives along local surface tangents, typically guided by estimated normals. However, coverage gaps may still arise when the initialization is suboptimal or when point density fails to adequately represent fine-scale structures.
@@
2D Foundational Models. Promptable segmentation models such as SAM [22] provide strong per-object cues that have been successfully applied in both vision and medical reconstruction [37, 32, 27, 70, 49]. For 3D Gaussians, SAGA [6] extends SAM-like capabilities but remains limited when handling fine-scale objects. In parallel, semantic backbones such as DINOv2 [39] and DINOv3 [50] provide dense, transferable feature embeddings that capture both local detail and global context. In particular, DINOv3 offers semantically consistent representations with strong invariance to viewpoint and appearance changes, making it especially well-suited for embedding-level supervision. This motivates our use of a DINOv3-based loss to complement pixel-domain objectives and enforce semantic alignment in the Gaussian Splatting pipeline.
@@
Point-based representations. Point clouds are flexible, permutation-invariant 3D carriers for reconstruction [64, 12, 34] and for tasks like 2D projection/detection [8] and point-level segmentation [42]. Point-based rendering learns to rasterize point features to images [43, 71], while self-supervised objectives densify or regularize point sets [56]. Broader advances span reconstruction [66, 41, 68, 10, 58, 51, 40], completion [60, 57, 69, 65], and point operators [52, 31, 16].
@@ 3.1 Convex Outlier Removal
To overcome these challenges, we employed a convex hull–based outlier removal method, introduced in [53], which leverages geometric properties to more reliably filter noise while preserving the integrity of the underlying structure. Unlike previous approaches, this method does not rely on local neighborhoods; instead, it evaluates each point independently and removes it if its distance to the convex hull exceeds a threshold.
@@ 3.2 Camera Clustering and Image Selection
Large multi-view datasets often contain hundreds of images, making segmentation across all views computationally prohibitive. Drawing inspiration from camera-parameter–driven view-reduction and calibration strategies [33, 21, 59, 26, 54], we propose a clustering-based approach to select a compact, geometry-aware subset of views.
@@
This approach preserves both spatial coverage and viewpoint diversity while substantially reducing the computational burden by limiting the number of images processed during segmentation. We apply SAM2 large model  [45] to the selected representative views, generating segmentation masks that include binary regions, bounding boxes, and area measurements. These masks are subsequently used to assign stable segment identifiers to 3D points for downstream processing tasks.
@@ 3.4 Mask-Area-Guided Point Cloud Enhancement
Previous research on point cloud upsampling has demonstrated that increasing point density significantly improves reconstruction quality [63, 24, 25, 46]. However, uniform densification fails to account for varying importance and visibility across different semantic regions within the scene. Our approach leverages segmentation mask areas from representative views to guide targeted point augmentation. Rather than applying uniform thresholds, we utilize mask area information to estimate appropriate point density for each segment, ensuring visually prominent regions receive adequate geometric representation. For each segment $s_{i}$ with mask area $A_{i}$, we determine the target number of points based on the square root relationship between area and required sampling density:
@@ 3.5 Feature Embedding Loss
We build upon recent advances in semantic feature–based supervision [15] by introducing a loss term grounded in DINOv3 feature embeddings [50]. Traditional photometric losses such as $\ell_{1}$ or SSIM capture only low-level pixel correspondences and thus fail to encode high-level semantic cues. To address this limitation, we incorporate a DINO-based supervision signal that enforces semantic consistency between ground-truth and rendered images. Prior works typically apply magnitude-based objectives such as $\ell_{1}$ distance directly on feature vectors, which conflates intensity differences with semantic correspondence. In contrast, we propose a cosine similarity formulation that emphasizes the angular relationship between embeddings, thereby isolating semantic correspondence while remaining invariant to feature magnitude scaling. Formally, given embeddings $f_{\text{gt}},f_{r}\in\mathbb{R}^{d}$ from ground-truth image $I_{\text{gt}}$ and rendered image $I_{r}$, we define
@@
This formulation integrates seamlessly into the 2D Gaussian Splatting framework [18]. The additional DINO term complements standard perceptual losses by aligning rendered views with the structural and semantic characteristics of the reference images. Consequently, each Gaussian primitive is optimized not only for photometric accuracy but also for semantic coherence, yielding reconstructions that better reflect the underlying scene semantics.
@@ 4 Experiments
We now present experiments for our GauSSmart approach, evaluating both appearance and geometry reconstruction across the three benchmark datasets used in this study. Specifically, we conduct experiments on DTU [19], Mip-NeRF 360 [3], and Tanks and Temples [23]. Our dataset preparation follows the same strategy introduced in the 2DGS paper [18].
@@ 4.1 Implementation Details
Loss Implementation. We employ DINOv3 with a Vision Transformer base architecture and $16\times 16$ patch size (ViT-B/16) as our feature extractor, selected after systematic evaluation across different DINOv3 variants on the Tanks and Temples dataset. The DINO loss weight is set to $\lambda=0.05$, determined through empirical evaluation over the range $[0.01,0.9]$ and guided by the weighting strategies described in [18].
@@
Outlier Removal. We adopt the convex-hull–based outlier removal procedure described in [53], applying it during both initialization and refinement to maintain geometric consistency across viewpoints.
@@ 5 Results
Our Gaussmart implementation demonstrates competitive performance across benchmark datasets. Table 1 reports a quantitative comparison against 3D Gaussian Splatting (3DGS)[20] and 2D Gaussian Splatting (2DGS)[18]. In terms of PSNR, our method outperforms 2DGS in four of six scenes, indicating reduced noise and improved fidelity. While 2DGS falls short of 3DGS on the Tanks and Temples dataset, the incorporation of our proposed loss and point cloud upsampling narrows this gap.
@@
In Fig. 6, our results are comparable to those of 3DGS, while exhibiting noticeably less noise than 2DGS. This improvement stems from reducing noise in the underlying point cloud, which often introduces bubble-like artifacts and leads to noisy renderings. Our reconstructions also preserve colors more consistently, confirming that the DINO-based loss captures and retains high-level semantic features. For outdoor scenes, our approach outperforms 2DGS across multiple datasets, demonstrating generalizability beyond indoor environments. This is particularly evident in the treehill scene, where complex structures such as root systems and tree trunks benefit from our feature-aware loss.
@@ 6 Ablations
Table 5 presents the mean PSNR across ablation configurations on Tanks and Temples. Introducing the DINO loss at iteration 7000 achieves a mean PSNR of 23.64, while activation at iteration 3000 results in slightly lower performance (23.61). Initializing the DINO loss from iteration 0 yields optimal performance, demonstrating that continuous semantic guidance throughout training provides superior reconstruction quality.

### v3

@@ 1 Introduction
Three-dimensional (3D) reconstruction has emerged as a prominent area of research since the introduction of Neural Radiance Fields (NeRF) [mildenhall2020nerf]. More recently, Gaussian Splatting [kerbl2023] has advanced scene reconstruction by exploiting the mathematical properties of 3D Gaussians, which allow for both accurate modeling and efficient rendering. Nevertheless, these methods frequently rely on unstructured point clouds, where uneven density and photogrammetry artifacts often introduce noise, limiting the fidelity of the reconstructed geometry.
@@
Given the spurious conditions of the input point clouds, we first apply a convex outlier removal procedure to eliminate unnecessary points introduced by interpolation errors in COLMAP reconstructions. In addition, we present a framework that leverages 2D foundation models such as DINOv3 [simeoni2025dinov3] and SAM [kirillov2023segment] to guide point cloud densification. Specifically, we increase the density of underrepresented segments by sampling new points according to the covariance structure of the original points within each segment mask. Finally, we introduce a novel loss function built on DINOv3 embeddings, which is differentiable and seamlessly integrates into the Gaussian Splatting pipeline [huang2024gaussian], enabling improved semantic alignment during optimization.
@@ 2 Related Works
Outlier Removal. Outlier removal is a key pre-processing step for noisy point clouds, with two widely used methods being Statistical Outlier Removal (SOR) and Radius Outlier Removal (ROR), both popularized in the Point Cloud Library [Rusu2011PCL, PCL_SOR_ROR]. SOR computes the mean distance of each point to its $k$-nearest neighbors and rejects those deviating from the global mean–standard deviation range, while ROR removes points with fewer than a threshold number of neighbors within a fixed radius. These lightweight filters are usually combined with robust model-fitting methods such as RANSAC [Fischler1981RANSAC] or Trimmed ICP [Chetverikov2002TICP], surface-fitting and projection approaches like Point Set Surfaces [Alexa2003PSS], graph-based and variational denoisers [Zeng2019GraphDenoise, Bouaziz2013L0], and learning-based frameworks such as PointCleanNet [Rakotosaona2019PointCleanNet].
@@
Novel View Synthesis Neural Radiance Fields (NeRF)[mildenhall2020nerf] model view-dependent color and density using a multilayer perceptron (MLP) to synthesize novel views. Subsequent variants have improved sampling strategies, scalability, and geometric fidelity[barron2021mipnerf, Liu20neurips_sparse_nerf, Lindell20arxiv_AutoInt, barron2022mipnerf360, yariv2023bakedsdf, cao2024lightning, oechsle2021unisurf, wang2021neus, yariv2021volume, meng_2023_neat, garbin2021fastnerfhighfidelityneuralrendering, lindell2021autoint]. Despite these advances, conventional NeRF models remain computationally expensive and often fail to accurately reconstruct large-scale or texture-poor regions. More recently, scene representations based on projected Gaussian primitives have enabled fast, high-quality rendering while remaining scalable to large scenes [kerbl2023, huang2024gaussian, choi2024meshgsadaptivemeshalignedgaussian, Dalal_2024, guédon2023sugarsurfacealignedgaussiansplatting]. In particular, 2D Gaussian Splatting [huang2024gaussian] improves geometric alignment by orienting primitives along local surface tangents, typically guided by estimated normals. However, coverage gaps may still arise when the initialization is suboptimal or when point density fails to adequately represent fine-scale structures.
@@
2D Foundational Models. Promptable segmentation models such as SAM [kirillov2023segment] provide strong per-object cues that have been successfully applied in both vision and medical reconstruction [morshuis2024segmentationguidedmrireconstructionmeaningfully, Ma_2024, li2024anatomaskenhancingmedicalimage, zhou2025distillationlearningguidedimage, shang2024jointsegmentationimagereconstruction]. For 3D Gaussians, SAGA [cen2025segment3dgaussians] extends SAM-like capabilities but remains limited when handling fine-scale objects. In parallel, semantic backbones such as DINOv2 [oquab2023dinov2] and DINOv3 [simeoni2025dinov3] provide dense, transferable feature embeddings that capture both local detail and global context. In particular, DINOv3 offers semantically consistent representations with strong invariance to viewpoint and appearance changes, making it especially well-suited for embedding-level supervision. This motivates our use of a DINOv3-based loss to complement pixel-domain objectives and enforce semantic alignment in the Gaussian Splatting pipeline.
@@
Point-based representations. Point clouds are flexible, permutation-invariant 3D carriers for reconstruction [yu2024pointdreamer, deng2020cvxnet, melaskyriazi2023pc2projectionconditionedpointcloud] and for tasks like 2D projection/detection [6923303] and point-level segmentation [qi2017pointnetdeeplearningpoint]. Point-based rendering learns to rasterize point features to images [rakhimov2022npbgacceleratingneuralpointbased, zhu2024rpbgrobustneuralpointbased], while self-supervised objectives densify or regularize point sets [Xiao_2023]. Broader advances span reconstruction [yu2022pointbertpretraining3dpoint, pang2022maskedautoencoderspointcloud, zhang2022pointm2aemultiscalemaskedautoencoders, dai2021can3dfast3dmedical, xu2023stop3dtargetreconstruction, sun20233dfusionrealtime3dobject, pan2023multidimensionunifiedswintransformer], completion [yang2021sanetshuffleattentiondeep, xie2018learningdescriptornetworks3d, zhang2021viewguidedpointcloudcompletion, yu2021pointrdiversepointcloud], and point operators [thomas2019kpconvflexibledeformableconvolution, liu2019relationshapeconvolutionalneuralnetwork, Guo_2021].
@@ 3.1 Convex Outlier Removal
To overcome these challenges, we employed a convex hull–based outlier removal method, introduced in [valverde2025convex], which leverages geometric properties to more reliably filter noise while preserving the integrity of the underlying structure. Unlike previous approaches, this method does not rely on local neighborhoods; instead, it evaluates each point independently and removes it if its distance to the convex hull exceeds a threshold.
@@ 3.2 Camera Clustering and Image Selection
Large multi-view datasets often contain hundreds of images, making segmentation across all views computationally prohibitive. Drawing inspiration from camera-parameter–driven view-reduction and calibration strategies [mauro2013overlapping, khan2019fast, yan2024livelearncontinualaction, article, waleed2024cameracalibrationgeometricconstraints], we propose a clustering-based approach to select a compact, geometry-aware subset of views.
@@
This approach preserves both spatial coverage and viewpoint diversity while substantially reducing the computational burden by limiting the number of images processed during segmentation. We apply SAM2 large model  [ravi2024sam2segmentimages] to the selected representative views, generating segmentation masks that include binary regions, bounding boxes, and area measurements. These masks are subsequently used to assign stable segment identifiers to 3D points for downstream processing tasks.
@@ 3.4 Mask-Area-Guided Point Cloud Enhancement
Previous research on point cloud upsampling has demonstrated that increasing point density significantly improves reconstruction quality [yu2018punetpointcloudupsampling, li2019puganpointcloudupsampling, li2021pointcloudupsamplingdisentangled, rong2024repkpu]. However, uniform densification fails to account for varying importance and visibility across different semantic regions within the scene. Our approach leverages segmentation mask areas from representative views to guide targeted point augmentation. Rather than applying uniform thresholds, we utilize mask area information to estimate appropriate point density for each segment, ensuring visually prominent regions receive adequate geometric representation. For each segment $s_{i}$ with mask area $A_{i}$, we determine the target number of points based on the square root relationship between area and required sampling density:
@@ 3.5 Feature Embedding Loss
We build upon recent advances in semantic feature–based supervision [gong2025dinoslamdinoinformedrgbdslam] by introducing a loss term grounded in DINOv3 feature embeddings [simeoni2025dinov3]. Traditional photometric losses such as $\ell_{1}$ or SSIM capture only low-level pixel correspondences and thus fail to encode high-level semantic cues. To address this limitation, we incorporate a DINO-based supervision signal that enforces semantic consistency between ground-truth and rendered images. Prior works typically apply magnitude-based objectives such as $\ell_{1}$ distance directly on feature vectors, which conflates intensity differences with semantic correspondence. In contrast, we propose a cosine similarity formulation that emphasizes the angular relationship between embeddings, thereby isolating semantic correspondence while remaining invariant to feature magnitude scaling. Formally, given embeddings $f_{\text{gt}},f_{r}\in\mathbb{R}^{d}$ from ground-truth image $I_{\text{gt}}$ and rendered image $I_{r}$, we define
@@
This formulation integrates seamlessly into the 2D Gaussian Splatting framework [huang2024gaussian]. The additional DINO term complements standard perceptual losses by aligning rendered views with the structural and semantic characteristics of the reference images. Consequently, each Gaussian primitive is optimized not only for photometric accuracy but also for semantic coherence, yielding reconstructions that better reflect the underlying scene semantics.
@@ 4 Experiments
We now present experiments for our GauSSmart approach, evaluating both appearance and geometry reconstruction across the three benchmark datasets used in this study. Specifically, we conduct experiments on DTU [jensen2014large], Mip-NeRF 360 [barron2022mipnerf360], and Tanks and Temples [Knapitsch2017]. Our dataset preparation follows the same strategy introduced in the 2DGS paper [huang2024gaussian].
@@ 4.1 Implementation Details
Loss Implementation. We employ DINOv3 with a Vision Transformer base architecture and $16\times 16$ patch size (ViT-B/16) as our feature extractor, selected after systematic evaluation across different DINOv3 variants on the Tanks and Temples dataset. The DINO loss weight is set to $\lambda=0.05$, determined through empirical evaluation over the range $[0.01,0.9]$ and guided by the weighting strategies described in [huang2024gaussian].
@@
Outlier Removal. We adopt the convex-hull–based outlier removal procedure described in [valverde2025convex], applying it during both initialization and refinement to maintain geometric consistency across viewpoints.
@@ 5 Results
Our Gaussmart implementation demonstrates competitive performance across benchmark datasets. Table 1 reports a quantitative comparison against 3D Gaussian Splatting (3DGS)[kerbl2023] and 2D Gaussian Splatting (2DGS)[huang2024gaussian]. In terms of PSNR, our method outperforms 2DGS in four of six scenes, indicating reduced noise and improved fidelity. While 2DGS falls short of 3DGS on the Tanks and Temples dataset, the incorporation of our proposed loss and point cloud upsampling narrows this gap.
@@
In Fig. 6, our results are comparable to those of 3DGS, while exhibiting noticeably less noise than 2DGS. This improvement stems from reducing noise in the underlying point cloud, which often introduces bubble-like artifacts and leads to noisy renderings. Our reconstructions also preserve colors more consistently, confirming that the DINO-based loss captures and retains high-level semantic features. For outdoor scenes, our approach outperforms 2DGS across multiple cases, demonstrating strong generalization beyond indoor environments. This is particularly evident in the treehill scene, where complex structures such as root systems and tree trunks benefit from our feature-aware loss.
For indoor environments, GauSSmart consistently achieves superior performance across all three quantitative metrics, surpassing both 2DGS and 3DGS. This improvement highlights the robustness of our method under controlled lighting conditions, where detailed textures and diverse scene elements are more effectively reconstructed.
@@ 6 Ablations
Table 5 presents the mean PSNR across ablation configurations on Tanks and Temples.